SOURCES Exercise 14-12 effect from the preferable -interest method) face amount of the bonds and redeemed them s also being amortized on a straight-line besis over the 30 years. (Straight-line is not materally different in eiy each December 31. The discount on The bors are calable at 101 (Le..t 101% of face-mount), and on January 2, 2017, Kingbird called Si oraring iacome aes, compute the amount ef los, any,to be recognized by kingbire as aresuit of retirig the $1,950,000 f onds in 2017.(Roum the $1,050,000 of bonds in 2017. (Round answer to O decimal places Loss on redemption Prepare th, yurna·erery to record the redentor. (Round answers to。decimal places, g.鷑548. Ir no entry is required, select No Entry for the account titles and amount is entered. Do not indent manually,) Date Account Titles and Explanation Debit Credit January 2, 2017 Question Attempts: 0 of 3 used BAVE FOR LATER SURNET ANSWER Version 4.24.11.
Answer #1
Discount Amount = 1050000*2/100 =21000
Yearly amortisation = 21000/10 = 2100
The loss recognised:
Amount $
Bonds called & redeemed 1050000*1.01 1060500
Less:Bonds payable -1050000
Add:Discount not amortised 21000-(2100*5) 10500
Loss on redeemption 21000
Journal Entry:
Date Accounts Titles and explanation Debit Credit
2-Jan-17 Bonds payable 1050000
Loss on Redeemption 21000
Discount on Bond issue 10500
Cash 1060500
(being bonds valuiing $1050000 got redeemed @
103 with adjustment of discount not amortised)